uhfinn:

Find the equation of a circle with its center at (2, 7) and a radius of 5.

uhfinn:

(x + 2)2 + (y + 7)2 = 25 (x + 2)2 + (y - 7)2 = 25 (x - 2)2 + (y + 7)2 = 25 (x - 2)2 + (y - 7)2 = 25

uhfinn:

answer choices^

Nnesha:

\[\rm (x-h)^2+(y-k)^2=r^2\] equation of a circle where (h,k) is the center and "r" is the radius.

uhfinn:

thank you